# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ 4.2.3 Running Time
Implementing Algorithm 3 requires standard dictionary operations—namely, inserting a key-value pair, deleting a key, and checking for key membership. These operations are used in multiple parts of the algorithm: during the blanket sampling and padding steps (Lines 8 and 9), and when maintaining the sparse representations of $\vec{h}$, $\hat{h}$, and $\tilde{h}$. The total number of such operations is bounded by $O(n+k)$.
If we rely on a balanced binary search tree (e.g., a red-black tree (Guibas and Sedgewick, 1978)), each operation takes $O(\log(n+k))$ deterministic time, resulting in a total runtime of $O((n+k)\log(n+k))=O(n\log n)$, since $k\in O(n)$.
We show how to replace these operations with integer sorting using a van Emde Boas tree (van Emde Boas, 1975), which allows sorting $n$ integers from the domain $[d]$ in $O(n\log\log d)$ time.
First, we construct $\vec{h}$—represented as a sorted array of (element, frequency) pairs—from the dataset $\mathcal{X}=\{{x^{({1})},\ldots,x^{({n})}}\}$ in $O(n\log\log d)$ time. Next, $\hat{h}$ can be obtained by copying $\vec{h}$ and replacing each frequency with a noisy value. This step takes $O(n)$ time, not including the cost of noise generation, which is accounted for separately.
The blanket sampling step proceeds by generating random elements with replacement and sorting them to remove duplicates, taking $O(n\log\log d)$ time.
The padding step then reduces to merging two sorted arrays, which takes $O(n)$ time. Finally, $\tilde{h}$ is constructed directly from the sorted array, also in $O(n)$ time.
4.3 From Replacement to Add/Remove Neighboring Model
We study the differentially private sparse histogram problem under the add/remove neighboring model. Unlike the replacement model, where the number of participants $n$ can be made public, in the add/remove model $n$ must remain private, as neighboring datasets may differ in size.
Our solution builds on the recent framework of Ratliff and Vadhan (2025) for converting differentially private, time-oblivious algorithms (referred to as Joint Output/Timing private algorithms in their work) from the upper-bounded setting (see the discussion following Definition 2.2) to the unbounded setting, under the add/remove neighboring model. The original framework is developed in the RAM model, where each memory cell can store an arbitrarily large natural number.
Proposition 4.1 ((Ratliff and Vadhan, 2025)).
$\mathds{TV}({\mathcal{M}{({y})}},\,{\mathcal{M}^{\prime}{({y})}})\leq\beta_{1}.$ (12)
For clarity of presentation, we tailor the framework to the private sparse histogram problem in word-RAM model. The pseudocode is given in Algorithm 4, and the construction relies on two key components.
Algorithm 4 Applying Ratliff and Vadhan’s Framework
2: parameter $\tau\in\mathbb{N}_{+}$triangleright as defined in Equation 2
3: histogram $\mathcal{X}=\{{x^{({1})},\ldots,x^{({n})}}\}$
5:for $k\in\mathbb{N}_{+}$ do
8:     $t\leftarrow\min\{n,\hat{n}_{k}\}$
9:     $\tilde{n}\leftarrow\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({\hat{n}_{k},\varepsilon_{k},\beta_{k}})}}\left({t}\right)$triangleright $\tilde{n}\approx\text{clamp}{[{{t+\operatorname{\mathbb{DL}ap}(e^{-\varepsilon_{k}})},{0},{\hat{n}_{k}}}]}$
10:     if $\tilde{n}<\frac{1}{2}\cdot\hat{n}_{k}$ then
11:         Truncate $\mathcal{X}$ to its first $\hat{n}_{k}$ entries if $n>\hat{n}_{k}$
12:         Construct $\vec{h}$ based on $\mathcal{X}$
13:         return $\mathcal{M}_{\textsc{\scriptsize Hist}}(d,\hat{n}_{k},k,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma},\vec{h})$triangleright Algorithm 3
Construction.
First, suppose we have an estimate $\hat{n}$ such that $n\leq\hat{n}\leq c\cdot n$ for some constant $c$. We run Algorithm 3 with $n$ replaced by $\hat{n}$, keeping all other parameters fixed. This modifies the algorithm in two ways: (1) noise is now sampled over the range $\mathbb{N}_{\leq{\hat{n}}}$, which preserves the original privacy guarantees and tail bounds; and (2) more items are sampled in line 8, which intuitively enhances the privacy of the selected item set $I$. This intuition is formally supported by revisiting the proof of Theorem 4.1, presented in Section 4.2. Although replacing $n$ with $\hat{n}$ increases the running time, the asymptotic complexity remains the same, since $\hat{n}\in O(n)$.
Second, Ratliff and Vadhan (2025) show that, for given $\beta_{1}\in(0,1)$, Algorithm 4 is an $\varepsilon_{1}$-DP time-oblivious algorithm that computes an estimate $\hat{n}$ satisfying $n\leq\hat{n}\leq c\cdot n$ with probability at least $1-\beta_{1}$. The algorithm constructs a sequence of candidate values $\hat{n}_{k}$, starting from a small initial guess and approximately doubling at each step, until it finds some $\hat{n}_{k}\geq n$. At each iteration, it spends privacy budget $\varepsilon_{1}/2^{k}$ to compare $\hat{n}_{k}$ to $n$: it proceeds with probability $1-\beta_{1}/2^{k}$ if $n\geq\hat{n}_{k}$, and halts with the same probability if $n\leq\hat{n}_{k}/2$. The privacy budget and failure probability are geometrically decayed to ensure that their total sums remain bounded by the overall privacy and failure parameters.
There is one caveat: the iteration may terminate with some $\hat{n}_{k}<n$, though this occurs with low probability. If we then run Algorithm 3 on a histogram constructed from the full dataset $\mathcal{X}$, using $\hat{n}_{k}$ in place of $n=\lvert{\mathcal{X}}\rvert$, the algorithm may fail. In particular, the privacy blanket sampling step (Algorithm 3, line 8) requires $\hat{n}_{k}+k\geq\lvert{I_{1}}\rvert$, where $\lvert{I_{1}}\rvert$ can be as large as $n$. To address this, Algorithm 4 truncates $\mathcal{X}$ to its first $\hat{n}_{k}$ elements before constructing $\vec{h}$ and invoking Algorithm 3. This truncation preserves the sensitivity of neighboring datasets and thus maintains privacy, at the cost of a small utility loss (with low probability).
Privacy, Utility Guarantees, and Running Time.
The joint output/time privacy guarantees and the utility bounds follow directly from Proposition 3.1 and the properties of Algorithm 3, as formalized in Theorem 4.1.
Analyzing the running time of Algorithm 4 in the word-RAM model requires more care. When $k\in O(\log n)$, the input parameters and internal variables such as $\varepsilon_{k}$, $\beta_{k}$, and $\hat{n}_{k}$ all fit within a constant number of words. If $\mathcal{M}_{\textsc{\scriptsize Hist}}$ (Algorithm 3) is not invoked, each iteration runs in $\tilde{O}(1/\varepsilon)$ deterministic time, and there are at most $O(\log n)$ such iterations.
When the algorithm terminates, it satisfies $\Pr[n\leq\hat{n}_{k}\leq c\cdot n]\geq 1-\beta_{1}$ for some constant $c$. Thus, with probability at least $1-\beta_{1}$, the cost of invoking $\mathcal{M}_{\textsc{\scriptsize Hist}}$ is $O(n\ln\ln d)$.
For larger $k\in\mathbb{N}$ such that $\hat{n}_{k}>c\cdot n$, the variables $\varepsilon_{k}$, $\beta_{k}$, and $\hat{n}_{k}$ may exceed word size and require $O(k)$ bits to represent. This increases the cost of both $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({\hat{n}_{k},\varepsilon_{k},\beta_{k}})}}\left({t}\right)$ and $\mathcal{M}_{\textsc{\scriptsize Hist}}$ by a factor of $O(k)$. However, since the probability of reaching iteration $k$ decays faster than geometrically once $\hat{n}_{k}>c\cdot n$, the overall expected running time remains bounded by $O(n\ln\ln d)$.
5 Efficient Private Noise Samplers
In this section, we present the noise samplers used in our private sparse histogram algorithm from Section 4. Our main result (Theorem 5.2) is an efficient, deterministic-time sampler that generates discrete noise over a finite interval, while achieving the same privacy guarantees and nearly matching the tail bounds of the discrete Laplace distribution.
This deterministic-time sampler belongs to a broader class known as time-oblivious samplers, introduced by Dov, David, Naor, and Tzalik (2023) and formally defined below. These samplers ensure that the runtime of generating each sample reveals little about the sampled value. This requirement is stronger than what our sparse histogram algorithm strictly needs—namely, that the total runtime of sampling all required noise values reveals only a controlled amount of information about the input histogram. Nonetheless, achieving per-sample time-obliviousness is desirable, as it enables safe use in scenarios where individual noise generation may be subject to timing side channels, such as in secure multiparty computation.
Our sampler (Theorem 5.2) is constructed in two steps: we first design a time-oblivious sampler (Theorem 5.3) whose output distribution closely approximates the discrete Laplace, and then apply the purification technique (Proposition 3.2) to convert its approximate DP guarantee to a pure one.
At the core of our construction is a simple, general framework for designing time-oblivious samplers for approximating arbitrary discrete distributions (Theorem 5.4), based on the Alias method (Algorithm 1). This framework improves upon a result by Dov et al. (2023), supports a wide range of differentially private noise distributions, and is of independent interest.
Time Oblivious Sampler.
Definition 5.1 ($(\varepsilon,\delta)$-Approximate Time-Oblivious Sampler (Dov et al., 2023)).
$e^{-\varepsilon}{({\mu{({E})}-\delta})}\leq\Pr_{\vec{s}\sim\operatorname{\mathbb{U}niform}\left({\{{0,1}\}^{\mathbb{N}}}\right)}\left[\mathcal{A}_{\mu}{({\vec{s}})}\in E\mid T_{{\mathcal{A}_{\mu}{({\vec{s}})}}}\in T\right]\leq e^{\varepsilon}\cdot\mu{({E})}+\delta,\quad\forall E\in\mathcal{F},$ (13)
When $\varepsilon=\delta=0$, the algorithm $\mathcal{A}_{\mu}$ is referred to as a (fully) time-oblivious generating algorithm for $\mu$ (Dov et al., 2023). However, the class of distributions that admit such samplers is quite limited: $\mu$ must have finite support, and all of its probabilities must be rational (Dov et al., 2023). This motivates the need for approximate relaxations in Definition 5.1.
5.1 Purified Approximate Discrete Laplace Sampler
The main result of this section is stated below.
Theorem 5.2 (Purified Approximate Discrete Laplace Sampler).
  • $\triangleright$
  • $\triangleright$
  • $\triangleright$
  • $\triangleright$
The first two properties in Theorem 5.2 state that $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma})}}$ matches the privacy guarantee and closely approximates the tail bound of the standard sampler $\mathcal{M}_{\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)}$, which creates shifted discrete Laplace noise: $\mathcal{M}_{\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)}(t)=t+\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ for any $t\in\mathbb{N}_{\leq{n}}$.
To prove the theorem, it suffices—by the purification technique (Proposition 3.2), which converts approximate DP to pure DP—to construct an efficient time-oblivious sampler whose output distribution is close to the discrete Laplace distribution.
Theorem 5.3 (Approximate Discrete Laplace Sampler).
  • $\triangleright$
  • $\triangleright$
Recall that Theorem 1.3 is a simplified version of Theorem 5.3. When $\varepsilon\doteq\frac{a_{\varepsilon}}{b_{\varepsilon}}$ and $\delta\doteq\frac{a_{\delta}}{b_{\delta}}$ for $a_{\varepsilon},b_{\varepsilon},a_{\delta},b_{\delta}\in\mathbb{N}_{+}$ that fit in a constant number of machine words, it holds that $\ln\frac{1}{\delta},\ln\frac{1}{\varepsilon}\in O(\omega)$. Hence, the memory usage becomes $O\left(\frac{1}{\varepsilon}+\ln\frac{1}{\delta}\right)$ words, and the sampling time becomes $O(1)$.
Before proving Theorem 5.3, we complete the proof of Theorem 5.2.
Proof of Theorem 5.2.
  1. 1.
    Initialize an instance of $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ stated in Theorem 5.3, with $\delta=\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\cdot\frac{\gamma}{1-\gamma}\cdot\frac{1}{1+n}$.
  2. 2.
  3. 3.
    Let $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma})}}\left({t}\right)=\mathcal{M}_{\operatorname{purify}}{({\mathcal{M}^{\prime},t,\gamma})}$, for each $t\in\mathbb{N}_{\leq{n}}$, where $\mathcal{M}_{\operatorname{purify}}$ is given in Algorithm 2.
Since $O{({\ln\frac{1}{\delta}})}=O{({\ln\frac{1+e^{-\varepsilon}}{1-e^{-\varepsilon}}+\ln\frac{1}{\gamma}+\ln n})}$, by Theorem 5.3, initializing $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ takes time $O{\big{(}{\frac{1}{\varepsilon}\cdot\textsc{poly}({\ln\frac{1}{\varepsilon},\ln n,\ln\frac{1}{\gamma}})}\big{)}}$ and memory usage $O{\big{(}{{\big{(}{\frac{1}{\varepsilon}+\ln\frac{1}{\gamma}+\ln n}\big{)}}\cdot{\big{(}{\ln\frac{1}{\gamma}+\ln n+\ln\frac{1}{\varepsilon}}\big{)}}}\big{)}}$. After initialization, by Propositions 3.2 and 5.3, $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma})}}\left({t}\right)$ can be eventuated in $O{\big{(}{\frac{1}{\omega}\cdot{\big{(}{\ln\frac{1}{\varepsilon}+\ln\frac{1}{\gamma}+\ln n}\big{)}}}\big{)}}$ time.
Privacy Guarantee.
Let $\mathcal{M}(t):\mathbb{N}_{\leq{n}}\rightarrow\mathbb{N}_{\leq{n}}$ be defined as $\mathcal{M}(t)\doteq\text{clamp}{[{{t+\operatorname{\mathbb{DL}ap}\left({e^{\varepsilon}}\right)},{0},{n}}]}$, which adds discrete Laplace noise to $t$ then clamp it to the range of $\mathbb{N}_{\leq{n}}$. It satisfies $\varepsilon$-DP: for each $t\in[n]$, $\mathcal{M}{({t-1})}$ and $\mathcal{M}{({t})}$ are $(\varepsilon,0)$-indistinguishable. By data processing inequality (Proposition A.1), we have
Utility Guarantee.
Since $\mathds{TV}({X},\,{\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)})\leq\delta$, for each $r\in\mathbb{N}^{+}$, based on the tail bound of discrete Laplace distribution (A.2),
Therefore,
5.2 Approximate Discrete Laplace Sampler
In this subsection, we prove Theorem 5.3. The proof proceeds in three steps.
First, we develop a general framework for constructing approximate time-oblivious samplers for arbitrary discrete distributions, not just the discrete Laplace. Our design is a direct instantiation of the Alias method (Algorithm 1), implemented truncated support and fixed point representation of probabilities.
Second, we instantiate this framework for the discrete Laplace distribution, obtaining a sampler that satisfies all the guarantees in Theorem 5.3 except for the memory usage.
Finally, to reduce space complexity, we reduce discrete Laplace sampling to geometric sampling. We leverage the fact that the geometric distribution has identical conditional distributions over intervals of equal length, which allows us to decompose a geometric sample into the sum of two smaller geometric components. Each component can then be approximately sampled using our general framework over smaller supports, further reducing the overall space usage.
5.2.1 Approximate Time-Oblivious Sampler Framework
Theorem 5.4 ($(0,\delta)$-Approximate Time-Oblivious Sampler).
  • $\triangleright$
    The labeling oracle $\mathcal{O}_{\text{label}}$, which assigns labels to the elements in $\operatorname{C}_{\delta/2}$ from 00 to $m-1$. For each $i\in\mathbb{N}_{<{m}}$, $\mathcal{O}_{\text{label}}(i)$ returns the corresponding element.
  • $\triangleright$
  • $\triangleright$
    It uses $2^{{\lceil{\log_{2}\lvert{\operatorname{C}_{\delta/2}}\rvert}\rceil}}\cdot\big{(}{\lceil{\log_{2}\frac{2}{\delta}}\rceil}+{\lceil{\log_{2}\lvert{\operatorname{C}_{\delta/2}}\rvert}\rceil}\big{)}$ bits of memory, which is in $O{({\lvert{\operatorname{C}_{\delta/2}}\rvert\cdot{({\ln\frac{1}{\delta}+\ln\lvert{\operatorname{C}_{\delta/2}}\rvert})}})}$.
  • $\triangleright$
Remark.
The detailed proof of Theorem 5.4 is deferred to Appendix E. Here, we present a proof sketch.
Proof Sketch for Theorem 5.4.
The pseudocode for the framework is presented in Algorithm 5. Let $m=\lvert{\operatorname{C}_{\delta/2}}\rvert$ and $\ell={\lceil{\log_{2}{({2/\delta})}+\log_{2}m}\rceil}$. We begin by retrieving the probabilities of elements in $\operatorname{C}_{\delta/2}$ to $\ell$ bits of binary precision via the oracle $\mathcal{O}_{\text{bin}}$.
The resulting truncated distribution has total variation distance at most $\delta$ from the original distribution $\mu$: the total mass outside $\operatorname{C}_{\delta/2}$ is at most $\delta/2$, and rounding errors within $\operatorname{C}_{\delta/2}$ contribute at most an additional $\delta/2$ by the union bound.
We then apply the Alias method (Algorithm 1) to sample exactly from this truncated distribution, without relying on the assumptions in Proposition 3.1—namely, that real numbers can be stored with infinite precision and that Bernoulli sampling can be performed with infinitely precise parameters.
The analysis of the space usage requires a careful examination of the initialization phase of the Alias method and is deferred to the full proof. ∎
Algorithm 5 Finite Alias Method ${\mathcal{M}_{\mathbb{FA}lias}}$
1: Procedure: Initialization
2: Input: distribution $\mu$ with discrete support, $\delta\in{({0,1})}$
3:$m\leftarrow\lvert{\operatorname{C}_{\delta/2}}\rvert$, $\ell\leftarrow{\lceil{\log_{2}{({2/\delta})}+\log_{2}m}\rceil}$
4:$\forall\,i\in\mathbb{N}_{<{m}}\,$: $x_{i}\leftarrow\mathcal{O}_{\text{label}}{({i})}$
5:$\forall\,i\in\mathbb{N}_{<{m}}\,$: $p_{i}\leftarrow\mathcal{O}_{\text{bin}}{({\mu{({x_{i}})},\ell})}$
6:${\mathcal{M}_{\mathbb{A}lias}.\textsc{\small Initialization}\left({m,p_{0},\ldots,p_{m-1}}\right)}$
7: Procedure: Sample
8: Output: Random variable $Z\in\operatorname{C}_{\delta/2}$
9:$Z\leftarrow{\mathcal{M}_{\mathbb{A}lias}.\textsc{\small Sample}\left({}\right)}$
10:return $\mathcal{O}_{\text{label}}{({Z})}$
5.2.2 Discrete Laplace Distribution
To apply Theorem 5.4 directly to the discrete Laplace distribution $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$, we identify the core support set $\operatorname{C}_{\delta/2}$ and specify the oracles $\mathcal{O}_{\text{label}}$ and $\mathcal{O}_{\text{bin}}$.
Core Support Set $\operatorname{C}_{\delta/2}$.
Due to the symmetric decay of the distribution’s probability mass from the center, the tail bound for the discrete Laplace distribution (A.2) implies that $\operatorname{C}_{\delta/2}=[{-L}\,.\,.\,{R}]$, where $L=R={\lceil{\frac{1}{\varepsilon}\ln\frac{4}{(1+e^{-\varepsilon})\delta}}\rceil}$.
Labeling Oracle $\mathcal{O}_{\text{label}}$.
Probability Oracle $\mathcal{O}_{\text{bin}}$.
Lemma 5.5 (Binary Expansion of Discrete Laplace Probability).
By Lemma 5.5, given $\ell={\lceil{\log_{2}{\big{(}{\frac{2}{\delta}}\big{)}}+\log_{2}\lvert{\operatorname{C}_{\delta/2}}\rvert}\rceil}$ and $t\in\operatorname{C}_{\delta/2}$, the $\ell$-bit binary expansion $\mathbb{B}inary\left({\Pr\left[\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)=t\right]},{\ell}\right)$ can be computed in time $O{({\textsc{poly}({\ln t,\ell})})}=O\big{(}\textsc{poly}(\ln\frac{1}{\varepsilon},\ln\frac{1}{\delta})\big{)}$. Consequently, $\mathcal{O}_{\text{bin}}$ can be implemented with total time $O{\big{(}{\frac{1}{\varepsilon}\cdot\ln\frac{1}{\delta}\cdot\textsc{poly}({\ln\frac{1}{\varepsilon},\ln\frac{1}{\delta}})}\big{)}}=O\big{(}\frac{1}{\varepsilon}\cdot\textsc{poly}(\ln\frac{1}{\varepsilon},\ln\frac{1}{\delta})\big{)}$, and requires an array of $O{\big{(}{\frac{1}{\varepsilon}\cdot\ln\frac{1}{\delta}\cdot\ell}\big{)}}=O{\big{(}{\frac{1}{\varepsilon}\cdot\ln\frac{1}{\delta}\cdot{\big{(}{\ln\frac{1}{\delta}+\ln\frac{1}{\varepsilon}}\big{)}}}\big{)}}$ bits to store the computed values.
Corollary 5.6 (Time-Oblivious Sampler for Discrete Laplace Distribution).
  • $\triangleright$
  • $\triangleright$
5.2.3 Memory-Efficient Construction
In this subsection, for $\varepsilon\in{({0,1})}$, we show how to reduce the $O{\big{(}{\frac{1}{\varepsilon}\cdot\ln\frac{1}{\delta}\cdot{\big{(}{\ln\frac{1}{\delta}+\ln\frac{1}{\varepsilon}}\big{)}}}\big{)}}$ bit memory usage in Corollary 5.6 to $O{\big{(}{{\big{(}{\frac{1}{\varepsilon}+\ln\frac{1}{\delta}}\big{)}}\cdot{\big{(}{\ln\frac{1}{\delta}+\ln\frac{1}{\varepsilon}}\big{)}}}\big{)}}$ bits, as stated in Theorem 5.3, at the cost of using slightly more random bits per sample, to complete the proof of Theorem 5.3.
The sampler construction is provided in Algorithm 6. It relies on two key decompositions. The first expresses a discrete Laplace random variable as a function of three simpler random variables: a Bernoulli, a uniform over $\{-1,1\}$, and a geometric. The Bernoulli and uniform variables can be (approximately) sampled directly.
The second decomposition breaks the geometric random variable into a composition of two simpler geometric variables, which are then sampled by our general framework based on the Alias method (Algorithm 5).
Algorithm 6 Discrete Laplace Noise ${\mathcal{M}_{\mathbb{DL}ap}}$
1: Procedure: Initialization
3:$r\leftarrow 2^{{\lceil{\log_{2}\frac{1}{\varepsilon}}\rceil}}$
4:${\mathcal{M}_{\mathbb{FA}lias}}^{\downarrow}\leftarrow{\mathcal{M}_{\mathbb{FA}lias}.\textsc{\small Initialization}}{({\operatorname{\mathbb{G}eo}{({e^{-\varepsilon},\mathbb{N}_{<{r}}})},\delta/3})}$
5:${\mathcal{M}_{\mathbb{FA}lias}}^{\uparrow}\leftarrow{\mathcal{M}_{\mathbb{FA}lias}.\textsc{\small Initialization}}{({\operatorname{\mathbb{G}eo}{({e^{-r\cdot\varepsilon}})},\delta/3})}$
6:$q_{\text{center}}^{\prime}\leftarrow\mathbb{B}inary\left({q_{\text{center}}=\Pr\left[\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)=0\right]},{\ell=\log_{2}\frac{3}{\delta}}\right)$
7: Procedure: Sample
9:$B\leftarrow\operatorname{\mathbb{B}ernoulli}\left({q_{\text{center}}^{\prime}}\right)$
10:$S\leftarrow\operatorname{\mathbb{U}niform}\left({\{{-1,1}\}}\right)$
11:$X\leftarrow{\mathcal{M}_{\mathbb{FA}lias}}^{\uparrow}.{\textsc{\small Sample}}{({})}$
12:$Y\leftarrow{\mathcal{M}_{\mathbb{FA}lias}}^{\downarrow}.{\textsc{\small Sample}}{({})}$
Decomposition of the Discrete Laplace Distribution.
The discrete Laplace distribution $\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)$ is defined by $\Pr\left[\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)=t\right]\propto e^{-\lvert{t}\rvert\cdot\varepsilon}$ for each $t\in\mathbb{Z}$. This distribution is symmetric about zero, and conditioned on $t>0$ or $t<0$, it reduces to a geometric distribution. This observation motivates the following sampling procedure: first, sample a Bernoulli random variable $B\sim\operatorname{\mathbb{B}ernoulli}\left({q_{\text{center}}}\right)$, where $q_{\text{center}}\doteq\Pr\left[\operatorname{\mathbb{DL}ap}\left({e^{-\varepsilon}}\right)=0\right]$, to decide whether to return 00; if not, sample a sign $S\sim\operatorname{\mathbb{U}niform}\left({\{{-1,1}\}}\right)$ to determine the output’s sign, and a geometric variable $X\sim\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}}\right)$ to determine the magnitude. This decomposition is formalized in 5.7, with a full proof in Appendix E.
Fact 5.7 (Discrete Laplace Decomposition).
One caveat is that $q_{\text{center}}$ is a real number, and thus cannot be computed exactly or sampled from $\operatorname{\mathbb{B}ernoulli}\left({q_{\text{center}}}\right)$ without approximation. However, as in Lemma 5.5, standard numerical techniques from Brent and Zimmermann (2010) allow us to compute an $\ell$-bit binary approximation $q_{\text{center}}^{\prime}=\mathbb{B}inary\left({q_{\text{center}}},{\ell}\right)$ in $\tilde{O}(\ell)$ time. As long as $\ell\in\Omega{({\log\frac{1}{\delta}})}$, we can ensure that $\mathds{TV}({\operatorname{\mathbb{B}ernoulli}\left({q_{\text{center}}}\right)},\,{\operatorname{\mathbb{B}ernoulli}\left({q_{\text{center}}^{\prime}}\right)})\in O(\delta)$. It remains to discuss how to approximately sample from $\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}}\right)$ within total variation distance $O(\delta)$.
Decomposition of the Geometric Distribution.
Applying our general framework (Theorem 5.4, Algorithm 5) to approximately sample from $\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}}\right)$ directly requires arrays whose lengths are proportional to the size of the core support set $\operatorname{C}_{\delta/2}=[{0}\,.\,.\,{{\lceil{\frac{1}{\varepsilon}\cdot\ln\frac{2}{\delta}}\rceil}}]$ (see Theorem 5.4 for the definition).
To reduce space usage, we cover the interval $[{0}\,.\,.\,{{\lceil{\frac{1}{\varepsilon}\cdot\ln\frac{2}{\delta}}\rceil}}]$ into smaller intervals of equal length: for $r\doteq 2^{{\lceil{\log_{2}\frac{1}{\varepsilon}}\rceil}}$, define $\mathcal{I}_{i}\doteq[{i\cdot r}\,.\,.\,{i\cdot r+r-1}]$ for each $i=0,\ldots,{\lceil{{\lceil{\frac{1}{\varepsilon}\cdot\ln\frac{2}{\delta}}\rceil}/r}\rceil}$. The key observation is that, conditioned on $Z$ falling into interval $\mathcal{I}_{i}$, the relative position $t$ within the interval follows identical distribution $\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}},{\mathbb{N}_{<{r}}}\right)$ (recall Definition 3.1): for $Z\sim\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}}\right)$,
This motivates a two-stage sampling procedure using independent geometric random variables: the first determines the interval index $i$, and the second determines the offset $t$ within $\mathcal{I}_{i}$. This decomposition is formalized in Lemma 5.8, with a complete proof given in Appendix E.
Lemma 5.8 (Geometric Decomposition).
Now we are ready to formally prove Theorem 5.3.
Proof of Theorem 5.3.
The precomputation cost is similar to that in Corollary 5.6. It therefore suffices to analyze the memory usage, sampling cost, and sample quality.
In Algorithm 6, the sampler ${\mathcal{M}_{\mathbb{FA}lias}}^{\downarrow}$ for $\operatorname{\mathbb{G}eo}\left({e^{-\varepsilon}},{\mathbb{N}_{<{r}}}\right)$, within total variation distance $\delta/3$, uses $O{\left({\frac{1}{\varepsilon}\cdot{\left({\ln\frac{1}{\delta}+\ln\frac{1}{\varepsilon}}\right)}}\right)}$ bits of memory, consumes $\log_{2}\frac{1}{\varepsilon}+\log_{2}\frac{1}{\delta}+O(1)$ unbiased random bits per sample, and runs in $O{\left({\frac{1}{\omega}\cdot\ln\frac{1}{\varepsilon}}\right)}$ time in the worst case.
[…]

### v2

@@ 4.2.3. Running Time
Implementing Algorithm˜2 requires standard dictionary operations— inserting a key–value pair, deleting a pair, updating the value associated with a key, and checking whether a key is present. These operations arise throughout the algorithm: during the histogram construction step (Line 4), during the blanket sampling and padding steps (Lines 8 and 9), and when reading or updating values stored in $\vec{h}$, $\hat{h}$, and $\tilde{h}$.
To support these operations, we use the hashing-based dictionary (Backyard Cuckoo Hashing) of Arbitman et al. (2010). Deleting a pair, updating a value, and checking whether a key is present all run in deterministic constant time per operation. Insertions require more care, as they may not always complete in constant time.
The key observation is that in Algorithm 2, there are $O(n+k)$ intersections in total. By Arbitman et al. (2010), for any constant $c$, the dictionary performs all these intersections in total time $O(n+k)$, with failure probability at most $(n+k)^{-c}$.
A natural approach is to enforce a deterministic time bound $T\in O(n+k)$, so that the combined dictionary operations exceed this bound with probability at most $(n+k)^{-c}$. However, this creates a subtle issue for differential privacy: two neighboring datasets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ may have different failure probabilities, potentially revealing information about their difference. A careful strategy is therefore required to control the impact of these failures on the overall privacy guarantee.
To handle potential failures, we adopt the following padding procedure, ensuring that neighboring datasets have similar failure probabilities with the hash dictionary:
  • $\circ$
    Flip a coin with probability $\frac{1}{e^{\varepsilon}-1}\cdot(n+k)^{-c}$. If heads, output a uniform random histogram.
  • $\circ$
    Otherwise, run Algorithm˜2. If all dictionary operations combined exceed time $T$, halt Algorithm˜2 and output a uniform random histogram $\tilde{h}$.
It remains to analyze its impact on the privacy guarantee. The output distribution of Algorithm 2 ($\mathcal{M}_{\textsc{\scriptsize Hist}}$) on input dataset $\mathcal{X}$, denoted $\bar{\mu}$, can be decomposed into two components: the distribution conditioned on the dictionary operations completing within the enforced time bound, denoted $\mu$, and a uniform distribution over all possible outputs, denoted $\nu$. There exist weights $w_{1},w_{2}\in[0,1]$ such that
$(1)\;\bar{\mu}=w_{1}\mu+w_{2}\nu,\;(2)\;w_{1}+w_{2}=1,\text{ and }\;(3)\;\frac{1}{e^{\varepsilon}-1}\cdot(n+k)^{-c}\leq w_{2}\leq\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\cdot(n+k)^{-c}.$
Similarly, for a neighboring dataset $\mathcal{X}^{\prime}$, let $\mu^{\prime}$ be the output distribution of $\mathcal{M}_{\textsc{\scriptsize Hist}}$ when the dictionary operations do not exceed the time bound. There exist weights $w_{1}^{\prime},w_{2}^{\prime}\in[0,1]$ such that
$(1)\bar{\mu}^{\prime}=w_{1}^{\prime}\mu^{\prime}+w_{2}^{\prime}\nu,\;(2)w_{1}^{\prime}+w_{2}^{\prime}=1,\text{ and }\;(3)\,\frac{1}{e^{\varepsilon}-1}\cdot(n+k)^{-c}\leq w_{2}^{\prime}\leq\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\cdot(n+k)^{-c}.$
Therefore, $\frac{w_{2}}{w_{2}^{\prime}}\leq e^{\varepsilon},\,\text{ and }\,\frac{w_{1}}{w_{1}^{\prime}}\leq\frac{1}{{1-\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\cdot(n+k)^{-c}}}.$ Further, based on the privacy analysis, $\mu$ and $\mu^{\prime}$ are $\varepsilon$-indistinguishable. For each measurable set $E$ in the range of $\mathcal{M}_{\textsc{\scriptsize Hist}}$, it holds that
$\frac{\bar{\mu}(E)}{\bar{\mu}^{\prime}(E)}$ $=\frac{w_{1}\,\mu(E)+w_{2}\,\nu(E)}{w_{1}^{\prime}\,\mu^{\prime}(E)+w_{2}^{\prime}\,\nu(E)}\leq\max\left\{{\frac{w_{1}\,\mu(E)}{w_{1}^{\prime}\,\mu^{\prime}(E)},\frac{w_{2}\,\nu(E)}{w_{2}^{\prime}\,\nu(E)}}\right\}\leq\frac{e^{\varepsilon}}{1-\frac{e^{\varepsilon}}{e^{\varepsilon}-1}\cdot(n+k)^{-c}}\leq e^{\varepsilon+\frac{2}{n^{2}}}.$
[…]